The Nubuck Shoe Company produces its famous shoe, the Divine Loafer that sells for $50 per pair. Operating income for NubuckWhat should Nubuck do? V has the highest operating income. the targeted increase in income of 25%* Requirement Evaluate each of the alternatives considered by Nubuck Shoes. Do any of the options meet or exceed Nubucks tar0 More Info 3. Replace a portion of its variable labor with an automated machining process. This would result in a 15% decreaData Table $ Sales revenue ($50 per pair) Variable cost ($20 per pair) Contribution margin 150,000 60,000 Fixed cost 90,000 3
Answer #1

Present structure:

No of units sold = $ 150,000/ $50 p.u. = 3000 unit

1.

1 2 3 4
Sales revenue (A) (No changes) 150,000 50$ * 3000* 1.3 = 195000 (50+15)*(3000*80%) = 156000 $ 50*(3000*150/100) = 225000
Variable Cost (B) 3000*(20*85%) = 51,000 20$ * 3000*1.3 = 78000 (20+8)*(3000*80%) = 67200 $ 20*(3000*150/100) = 90000
Contribution (C = A-B)            99,000 117000 88800 135000
Fixed Costs (D) 36000*110/100 = 39,600 36000+20000 = 56000 (no changes) 36000 36000*2 = 72000
Operating Income (C - D)            59,400 61000 52800 63000

2.

Operating Income (from above) 59400 61000 52800 63000
Existing profit (E) 54000 54000 54000 54000
Difference (F) 5400 7000 -1200 9000
Percentage of profit (E/F)*100 10.00% 12.96% -2.22% 16.67%
Met/exceeded? NO NO NO NO

Evaluation:

None of the alternatives meet criteria of 25% increase in profits. The best one is Alternative 4 yields 16.67%.

Hence, Nubuck should look into using combination of two alternatives: It can combine Alternative 4 ie., adding second manufacturing facility and also Alternative 2 ie spending on advertisement or Alternative 4 along with alternative 1 ie reducing variable costs p.u. these can increase profit by 25%.

Please comment in case of any query regarding the solution.